10. What is AG in kJ for the reaction: cua) 2 Br Br20) 2 Cl if Br]-0.100, DC11 0.50, and p Clz) 1.00 atm? The temperature is 25°C. Final Answer -48 J
Answer #1

Ans. Cl2(g) + 2 Br-(aq) = Br2(l) + 2 Cl-(aq)

Equilibrium constant of the reaction, K = ([Br2] [Cl-]2) / ([Cl2] [Br-]2)

The concentration of pure chemical species is taken to be unity.

So, K = [Cl-]2 / [Br-]2 = (0.50)2 / (0.10)2 = 25

Now,

At equilibrium,

Using the equation dG = dG0’ + RT lnK              - equation 1

            Where, dG = calculated/ experimental free energy change = ?

                                    dG0’ = standard/ theoretical free energy change                                                            R = 0.0083146 kJ mol-1K-1

                                    T = temperature in kelvin = (0C + 273.15) K

                                    K = equilibrium constant under given condition.

We have, dG0 for the reaction = - 56.35 kJ/mol

Putting the values in equation 1-

            dG = - 56.35 kJ mol-1+ (0.0083146 kJ mol-1 K-1) x 298.15 K x (ln 25)

            Or, dG= - 56.35 kJ mol-1 + 2.47899799 kJ mol-1 x (3.219)

            Or, dG= - 56.35 kJ mol-1 + 7.98 kJ mol-1

            Or, dG = - 48.37 kJ mol-1

Therefore, dG of the reaction = - 48.37 kJ mol-1

Note: The value of dG shall be in terms of kJ/ mol. Please use the dG0 mentioned in your textbook top get the dG value closer to -48 kJ/mol or round off -48.33 kJ/mol to -48 kJ/mol.